2.44gm of Benzoic acid is dissolver in benzene to  form 250 ml of soln with dimer structure what is formality 

Grade:11

1 Answers

Shahzada
5 years ago
F=noof moles of solute /one unit of solvent 
 M Mass of Benzoic acid =122 as formula
C6H5COOH
Find mole =2.44/122=0.02
That is dimer multiple by 2 
122×2=244. 0.25×2=0.5
F=0.02/244×0.5